# Reviewing Intelligent Cinematography: AI research for camera-based video production

### v1

@@ 3.2.2 Camera Calibration and Localization
The objective of camera calibration and localization (or pose estimation) is to map a 3-D world onto a 2-D image plane. This involves modelling camera intrinsic/inertial parameters and extrinsic/external parameters [295] using real image data. An illustration of the relevant parameters to be modelled is shown in Figure 12. The intrinsic parameters define the cameras model, focal length, and lens distortions. The extrinsic parameters model the camera transform matrix for each pose as well as the path for moving shots.
@@
Figure 12: Extrinsic parameters are shown in grey. Intrinsic parameters are shown in white.
@@
This is a popular problem that concerns the general use case of a single camera [198, 152, 208] as well as specific use cases, such as surgical monitoring [198, 175, 130]. Regarding IC, we find work supporting a number of applications including aerial photography [25, 195], photogrammetry [154, 59, 77], image-based 3-D reconstruction [47, 250, 275, 170] and underwater filming [230, 38, 155, 163, 299]. The two practical uses for this in cinematography are automated camera control and real 3-D reconstruction.
@@
For automated camera control, the research landscape leans towards controlling the extrinsic parameters [219, 208]. However, there exists work bridging extrinsic and intrinsic parametric control to satisfy the cinematographer. For example, Pueyo et al [195] present CineMPC which searches for an optimum trajectory (for a drone) and camera angle using a model predicted control (MPC) framework. MPC [33] achieves a process output such as a camera action by considering future time instances/horizons and minimising the cost of selecting different actions. CineMPC models a finite horizon which is continually displaced until all actions cease and constrains the objective function to be flexible to different camera configurations and visual aesthetics. The authors curate mathematical expressions to account for composition, depth of field and canonical shots. Thus, the subsequent cost functions can be considered naive approximations of canonical cinematographic style.
@@
While there exists a lot of work on camera calibration for automated control, the landscape with regards to 3-D reconstruction is more fertile. Discussed by Remondino and Fraser [208], the general approaches to calibration are Tsai [251], Heikkila and Silvén [95] and Zhang [294] which model distortion intrinsic parameters for a pinhole camera. Nevertheless, there is no universal or flexible automation scheme that achieves this [208]. This is why we still see splintered research on subsets paradigms such as Pitombeira and Mitishita [191] which focuses on calibrating the zoom and focus features for scenarios where modelling fixed focus and zoom lenses are infeasible. Or Jasińska et al [113], which looks at improving geometric stability of videos using SfM of multi-view stereo (MVS) images for photogrammetric reconstruction.
@@
Interestingly, there is work relating to INR and NeRF modelling which addresses the whole problem [300]. The classical model in Mildenhall et al [166] naively uses COLMAP [226] whereas recent work has focused on joint-optimization [279, 274, 267, 250]. While many of these papers improve upon pose estimation for a fixed pinhole camera model, there is work on evolving the lens model as well. For example, Xian et al [275] presents a formulation of a ResNet-based lens distortion model and robust patern-based calibration to provide a thin-lens model equitable for NeRF reconstruction as well as other vision based tasks.
@@ 3.2.3 Neural 3-D Representations of Dynamic Scenes
Method
Speed
Commercial GPU
Explicit
Ancestors
@@
K-Planes m NeRF-W, Instant-NGP, DVGO
HyperReel h TensorRF
@@
Table 2: Overview of Dynamic NeRFs: Speed measures the training time where D is days, h is hours and m is minutes. Commercial GPU indicates whether the authors tested with commercially available GPUs. Explicit indicates where an explicit representation was used to boost training and inference speed. Ancestors indicates the prior work influencing each method.
There are two traits that currently differentiate originality of research. The first is the proposal of new space-time representations. The second is the manipulation of representations to enhance the learning of temporal elements.
One of the earliest and easiest-to-understand proposals for a new space-time representation is D-NeRF [196]. This method models the deformation field $\Phi(\textbf{x},t)\rightarrow\Delta\textbf{x}$ where $\Delta\textbf{x}$ is the predicted positional change of a ray-sample relative to a canonical static field $\Phi(\textbf{x}+\Delta\textbf{x},t)\rightarrow(\textbf{c},\sigma)$. Relative to global space, this learns an SE(3) transformation. Learning a canonical static space is a robust way of ensuring volumetric consistency with time. Though intuitive, this lends itself to issues when a scene is not continuously in-frame. To counterbalance this, methods such as K-Planes decompose static and dynamic volumes representations. More specifically, K-Planes does this by projecting ray samples (containing $(\textbf{x},t)$) into 6 feature planes, three representing static space and three representing dynamic space. The inputs are normalized between $[0,N]$, projected onto the feature planes and bi-linearly interpolated among varying scales (i.e. coarse and fine features). To decode the features, attaining $(c,\sigma)$, element-wise multiplication is used to recover a final feature vector which is passed into a feature decoder (for explicit representation) or small MLP (for implicit representation). Overcoming similar problems with large and slow-to-train dynamic representation we see GS alternatives such as 4DGS [271] [^12] Pre-print available on arXiv, which use the same K-Plane decompisition, though use it to derive covariance and visual properties, rather than visual-only properties as with K-Planes.
Another interesting solution to the issue of volumetric consistency is key-frame interpolation of static fields. HyperReel exemplifies this by learning the displacement vectors and geometric primitives of a jointly learned key-framed static field. To learn many key-frame fields for a single video, HyperReel builds upon TensorRF due to its compact nature and fast learning ability. Furthermore, as this only provides a discrete set of time-dependant snapshots the authors propose modelling the velocity of volumes at each key-frame. Similarly to the principles of D-NeRF, this enhances the temporal quality of the radiance field representation.
Overall, methods are only modestly capable of modelling dynamic scenes with 6-DoF and usually require additional modification to handle non-generic scenes, such as forward-facing [10, 78]. Despite this, we still find work that tends to production specific needs such as editable NeRFs and GSs’ representations [136, 298, 109]. For cinematographers this means patience as we continue to witness increasing interest in this field.
@@ 3.2.4 Neural 3-D Representations of Humans
Modelling non-rigid or deformable geometry, for example a human, is a classical problem for graphics research. This topic is well aligned with research on dynamic NeRFs and broadens human-centered computer vision research to the 3-D case.
Similarly to the prior subsection, the two general problem formulations are multi-view and monocular scenes, which we use to distinguish current research objectives. We can additionally differentiate work by it’s reliance on generalising humans and their poses. HumanNeRF [296] exemplifies both of these points as a method that focuses on the multi-view paradigm while generating generalizable human poses. The authors propose using a NeRF to learn the skinner multi-person linear model (SMPL) [153], i.e. the geometry and appearance of an actor. These features are learnt and used to train a novel neural appearance blending field. Similarly to D-NeRF, the generalizable NeRF learns the canonical and deformation field of an actor by taking inputs of an SMPL skeleton and pixel aligned features and outputting the colour and density of a sample. The appearance blending field refines texture details by accounting for the colour of aligned features from neighbouring views. This model performs well, however it requires carefully placed cameras and struggles with new poses.
@@
In cinematographic practice, neural human modelling methods relives us of the accessibility constraints of motion capture (MoCap) suits [260, 261] - which are currently viewed as the gold standard MoCap systems. The authors of the neural motion (NeMo) model [261] exemplify this by testing their proposed framework on athletic actions, taken from the Penn Action Dataset [292]. Like MonoHuman and HumanNeRF, NeMo generalizes motions using multi-view video. Though notably this is achieved by inputting videos of the same action with varying scene conditions, such as different actors and lighting. To handle un-synchronized actions a phase network is introduced as time-based warping function to align poses (i.e. joint angles and translation of motion) in an action sequence. Furthermore, scene-specific feature embeddings handle the visual differences between scenes. Consequently, the phase and instance features are used as inputs to the NeMo network where the outputs define the joint angles and translation of motion - used to render an SMPL model.
Conclusively, this field of research is popular and highly relevant to IC. While it is currently limited by training and inference time [261, 286] there is potential for it use in applications like MoCap, which may be additionally beneficial for actor-based ICVFX.
@@ 3.2.5 Challenges and Future Work
With ICVFX, research is limited by access to expensive technologies such as cameras and LED volumes. Considering the novelty of this production format, we are certain to see more research emerge as more stages are built.
Aside from current research on improving the quality of stage lighting, there are other areas that we believe are worth pursuing. LED volume productions (or virtual stage production (VSP)) is heavy on pre-production planning, thus there is potential for supporting tasks like pre-production shot visualization, or 3-D stage design. This area of research branches well with 3-D capture and modelling research. Real-time 3-D view-synthesis (rendering) could also play a role in in-camera background projection for VSP. Though prior to this there are steps yet to be taken in providing high-resolution renders or real-time scene changes.
With the recent introduction of open-source pipelines for NeRF [243, 140] modifying and testing different NeRFs is trivial. However, there is a lack of understanding when it comes to the optimal strategy of capturing media for training NeRFs. Additionally, current methods for evaluating NeRFs using 2-D image-based metrics such as PSNR, SSIM and LPIPS are perhaps not completely reflective of the accuracy of the 3-D fields which we expect a NeRF network to learn [243, 83]. This is specifically true for the dynamic use-case. Therefore, we believe this field still has a considerable journey ahead before we can confidently select models for large scale commercial productions. It is also worth noting that automated 3-D capture technologies play a great role in the general worry over the ownership of an actors visual essence. In Section 4.1 we elaborate on the social responsibility of researchers as we find this ethical dilemma akin to issues previously faced with Deep Fake research.
@@ 3.3 Live Production
IC research for live production is conditional on the type of environment, types of actions expected to be displayed (e.g. routines or phases in a performance or sport), location of audience members and the focal point for the digital audience. Analysing the current play or performance can live selection of cameras and shots. Additionally, challenging environments and scenarios may lead to degraded visual quality, thus real time solutions are required for image correction. For example, image analysis from a game of football is challenged by noise arising from spectators or advertisement banners [187, 238, 270], while water-sports will face problems of distortion from partially underwater participants, light artifacts from surface reflection and noise from turbulent water [270, 104]. In this section we link work concerning more research on object tracking and human pose estimation and specifically highlight cases that could inform live shot selection or real-time image correction methods.
@@ 3.3.1 Human Pose Estimation
Human pose estimation (HPE) is a popular paradigm in computer vision research [11, 223, 297]. We see cases of HPE used in cinematic scene analysis [265, 273] and also in live event broadcasting. For example, for sport events analysing a player’s pose may allow us to to forecast a series of entertaining events which could provoke particular cinematographic shots. For sport there is ultimately one major factor that divides research: whether or not general application is possible. This is an essential delineation to make prior to reviewing a methods [11] as solutions are heterogeneous [^13] E.g. anatomical poses will vary between sport.. Aside from this there are special cases where models take-on specific challenges, such as exploring HPE in 3D [265, 237], in a highly dynamic environments [98] or for team sports [27].
Hu [105] looks at real-time football player (single-target) posture estimation for live presentational analysis. To accomplish this, confidence weighted fusion and visual attention is used handle problems with colour camouflage and static foreground features, to first identify the target foreground features. Figure13 illustrates pixel-based joint verification for identifying key-point target features, using local binary similarity patterns (lbsp). Then a heat map is generated using a ResNet, a stacked hour glass network and deep-layer aggregation (DLA). The DLA collects features from CNN layers to determine the shared features, i.e. it aggregates features at each layer relating to a verified joint and tries to classify what it is (knee, elbow, etc.).For optimization, the model is updated with adaptive strategies relating to the confidence weight of each pixel and their corresponding weighted fusion sum - based on the joint classification. This methods shows significant improvement in joint and posture detection compared to other state-of-the-art methods [257, 190, 103, 184].
(a) Joint-verification using color features, $v_{t}$, and the local binary similarity pattern, $lbsp_{t}$, of pixel $x$ at frame $t$. $R_{color}$ and $R_{lbsp}$ are predefined distance thresholds and $\{v_{1},v_{2},...,v_{n}\}$ and $\{lbsp_{1},lbsp_{2},...,lbsp_{m}\}$ are a set of known color and binary similarity features. The distance function between the pixel’s colour feature, $lbsp$ feature and features in their respective sets, is evaluated w.r.t the thresholds and the binary result is fused using a logical AND operation to determine where a pixel belongs to a joint.
@@
Figure 13: Per-pixel binary joint verification used for identifying pixels for joint classification.
Differently, Bridgeman et al [27] looks at tracking and analysing multiple targets simultaneously. This method builds upon existing research by Cao et al [36]. The goal is to construct a 3D (skeletal) pose model using multiple 2D views, while tackling prior issues of long processing time and reliance on appearance-based queues for initiating feature detection. This is achieved by presenting a greedy approach to multi-view detection. The approach has three steps: (1) 2D Pose Error Correction is done by flipping incorrectly oriented body parts and dividing by body part, (2) Per-frame 2D pose association determines a consistent label for each body part across multiple views, found greedily by selecting the best pose from a weighted rank, and (3) 3D skeleton tracking uses the 2D labelled poses to generate a 3D skeleton for each individual. The novelty in this paper is it’s use of multi-view HPE to verify joint placements in 3D.

### v3

@@ 3.2.2 Camera Calibration and Localization
The objective of camera calibration and localization (or pose estimation) is to map a 3-D world onto a 2-D image plane. This involves modeling camera intrinsic/inertial parameters and extrinsic/external parameters [308] using real image data. An illustration of the relevant parameters to be modeled is shown in Figure 12. The intrinsic parameters define the cameras model, focal length, and lens distortions. The extrinsic parameters model the camera transform matrix for each pose as well as the path for moving shots.
@@
Figure 12: Visualizing the parameters involved in camera calibration: Extrinsic parameters are shown in grey. Intrinsic parameters are shown in white.
@@
This is a popular problem that concerns the general use case of a single camera [209, 162, 219] as well as specific use cases, such as surgical monitoring [209, 185, 139]. Regarding IC, published work relates to several applications including aerial photography [25, 206], photogrammetry [164, 61, 82], image-based 3-D reconstruction [49, 262, 288, 180] and underwater filming [242, 39, 165, 173, 312]. The two practical uses for this in cinematography are automated camera control and real 3-D reconstruction.
@@
For automated camera control, the research landscape leans towards controlling the extrinsic parameters [231, 219]. However, there exists work bridging extrinsic and intrinsic parametric control to satisfy the cinematographer. For example, Pueyo et al [206] present CineMPC which searches for an optimum trajectory (for a drone) and camera angle using a model predicted control (MPC) framework. MPC [34] achieves a process output such as a camera action by considering future time instances/horizons and minimizing the cost of selecting different actions. CineMPC models a finite horizon which is continually displaced until all actions cease and constrains the objective function to be flexible to different camera configurations and visual aesthetics. The authors curate mathematical expressions to account for composition, depth of field and canonical shots. Thus, the subsequent cost functions can be considered naive approximations of canonical cinematographic style.
@@
While there exists a lot of work on camera calibration for automated control, the landscape with regard to 3-D reconstruction is more fertile. As discussed by Remondino and Fraser [219], the general approaches to calibration are Tsai [263], Heikkila and Silvén [101] and Zhang [307] which model distortion intrinsic parameters for a pinhole camera. Nevertheless, there is no universal or flexible automation scheme that achieves this [219]. This is why we still see splintered research on subsets paradigms such as Pitombeira and Mitishita [201] which focuses on calibrating the zoom and focus features for scenarios where modeling fixed focus and zoom lenses are infeasible. Or Jasińska et al [120], which looks at improving geometric stability of videos using SfM of multi-view stereo (MVS) images for photogrammetric reconstruction.
@@
Interestingly, there is work relating to INR and NeRF modeling which addresses the whole problem [313]. The classical model in Mildenhall et al [176] naively uses COLMAP [238] whereas recent work has focused on joint-optimization [292, 287, 279, 262]. While many of these papers improve upon pose estimation for a fixed pinhole camera model, there is work on evolving the lens model as well. For example, Xian et al [288] presents a formulation of a ResNet-based lens distortion model and robust patern-based calibration to provide a thin-lens model equitable for NeRF reconstruction as well as other vision based tasks.
@@ 3.2.3 Neural 3-D Representations of Dynamic Scenes
Method Training Speed Compact Explicit Ancestors
@@
K-Planes m/h NeRF-W, Instant-NGP, DVGO
HyperReel m/h TensorRF
@@
Table 2: Overview of Dynamic NeRFs: Training Speed measures the training time where D is days, h is hours and m is minutes. This may be scene-dependent so m/h indicates that training speed varies from minutes to hours. Compact indicates whether the models can be trained on a single (commercially available) GPU. Explicit indicates where an explicit representation was used to boost inference speed. Ancestors indicates the prior work influencing each method
There are two traits that currently differentiate research approaches. The first is the proposal of new space-time representations. The second is the manipulation of representations to enhance the learning of temporal elements.
One of the earliest and easiest-to-understand proposals for a new space-time representation is D-NeRF [207]. This method models the deformation (change in position) of volumes w.r.t time. Though intuitive, deformation models present issues when a scene/objects are not continuously in-frame. Instead, methods like K-Planes Fridovich-Keil et al [83] use low-rank decompositions (e.g., 4-D to 2-D projections) to consistently model scene topology even when objects go unseen for several frames. We discuss D-NeRF and K-Planes in more technical detail in Appendix A.2. To overcome issues with large, slow-to-train dynamic representations, Gaussian Splatting alternatives such as 4DGS [284] are preferable. These use the same K-Plane decomposition but apply it to predict both the visual and positional properties of Gaussian blobs, rather than only the visual properties as K-Planes does.
Another interesting solution to the issue of volumetric consistency is key-frame interpolation of static fields. HyperReel exemplifies this by learning the displacement vectors and geometric primitives of a jointly learned key-framed static field. To learn many key-frame fields for a single video, HyperReel builds upon TensorRF due to its compact nature and fast learning ability. Furthermore, as this only provides a discrete set of time-dependent snapshots the authors propose modeling the velocity of volumes at each key-frame. Similarly to the principles of D-NeRF, this enhances the temporal quality of the radiance field representation.
Overall, methods are only moderately capable of modeling dynamic scenes with 6-DoF and usually require additional modification to handle non-generic scenes, such as forward-facing [9, 83]. Despite this, we still find work that tends to production specific needs such as editable NeRFs and GSs’ representations [146, 310, 117]. For cinematographers this means being patient as we witness increasing interest and contributions in this field.
@@ 3.2.4 Neural 3-D Representations of Humans
modeling non-rigid or deformable geometry, for example a human, is a classical problem for graphics research. This topic is well aligned with research on dynamic NeRFs and broadens human-centered computer vision research to the 3-D case.
Similarly to the prior subsection, the two general problem formulations are multi-view and monocular scenes, which we use to distinguish current research objectives. We can additionally differentiate work by it’s reliance on generalizing humans and their poses. HumanNeRF [309] exemplifies both of these points as a method that focuses on the multi-view paradigm while generating generalizable human poses. The authors propose using a NeRF to learn the skinner multi-person linear model (SMPL) [163], i.e. the geometry and appearance of an actor. These features are learned and used to train a novel neural appearance blending field. Similarly to D-NeRF, the generalizable NeRF learns the canonical and deformation field of an actor by taking inputs of an SMPL skeleton and pixel aligned features and outputting the color and density of a sample. The appearance blending field refines texture details by accounting for the color of aligned features from neighbouring views. This model performs well, however it requires carefully placed cameras and struggles with new poses.
@@
In cinematographic practice, neural human modeling methods relive the accessibility constraints of motion capture (MoCap) suits [272, 273] - which are currently viewed as the gold standard for MoCap systems. The authors of the neural motion (NeMo) model [273] exemplify this by testing their proposed framework on athletic actions, taken from the Penn Action Dataset [305]. Like MonoHuman and HumanNeRF, NeMo generalizes motions using multi-view video. Though notably this is achieved by inputting videos of the same action with varying scene conditions, such as different actors and lighting. To handle un-synchronized actions a phase network is introduced as time-based warping function to align poses (i.e. joint angles and translation of motion) in an action sequence. Furthermore, scene-specific feature embeddings handle the visual differences between scenes. Consequently, the phase and instance features are used as inputs to the NeMo network where the outputs define the joint angles and translation of motion - used to render an SMPL model.
Conclusively, this field of research is popular and highly relevant to IC. While it is currently limited by training and inference time [273, 299] there is potential for its use in applications like MoCap, which may be additionally beneficial for actor-based ICVFX.
@@ 3.2.5 Challenges and Future Work
With ICVFX, research is limited by access to expensive technologies such as cameras and LED volumes. Considering the novelty of the LED volume production format, we are certain to see more research emerge as more stages are built. Hogg [110] echos this by presenting commentary from numerous production supervisors and specialists who demonstrate grand expectations for the technology.
Aside from current research on improving the quality of stage lighting, there are other areas offering potential. LED volume productions (or virtual stage production (VSP)) have become popular, relying heavily on pre-production planning; hence there is potential for supporting tasks like pre-production shot visualization, or 3-D stage design. This area of research aligns well with 3-D capture and modeling research. Real-time 3-D view-synthesis (rendering) could also play a role in in-camera background projection for VSP. This will involve resolving high-resolution real-time rendering problems associated with projecting large live backgrounds.
With the recent introduction of open-source pipelines for NeRF [255, 150], modifying and testing different NeRFs has become trivial. This has been showcased by popular online creatives such as Wren Weichman (a member the Corridor Crew YouTube channel) who demonstrates the use of NeRF for numerous CGI and VFX tasks like asset generation and background projection in a video titled “I try the tech that WILL replace CG one day” [^9] Accessible here: https://www.youtube.com/watch?v=YX5AoaWrowY. This could be elevated for use in an ICVFX pipeline by combining NeRF/GS methods that model backgrounds with real-time automated foreground segmentation techniques. The benefit to this over existing ICVFX solutions is that we no longer need to manually model realistic virtual background and can instead model digital twins of real environments, which is significantly less time consuming and costly. As a current example, the BBC have been investigating NeRFs for synthesizing new shots that are challenging to acquire manually [^10] Accessible here: https://www.bbc.co.uk/programmes/p0gz4rr0. They are able to achieve this in practice due to a well conceived, scene-specific method of capture. However, this demonstrates that there remains a lack of understanding regarding the optimal strategy for capturing media to train NeRFs. Additionally, current methods for evaluating NeRFs using 2-D image-based metrics such as PSNR, SSIM and LPIPS are perhaps not completely reflective of the accuracy of the 3-D fields which we expect a NeRF network to learn [255, 88]. This is specifically true for the dynamic use-case. Hence, this field still requires further work before we can confidently select models for large scale commercial productions. It is also worth noting that automated 3-D capture technologies impact on current concerns over the ownership of an actors visual essence. In Section 3.5 we elaborate on the social responsibility of researchers relating to this ethical dilemma.
@@ 3.3 Live Production
IC research for live production is conditional on the type of environment, types of actions expected to be displayed (e.g., routines or phases in a performance or sport), location of audience members and the focal point for the digital audience. Additionally, challenging environments and scenarios may lead to degraded visual quality, thus real time solutions are required for image correction or enhancement. For example, image analysis from a game of football is challenged by noise arising from spectators or advertisement banners [197, 250, 283], while water-sports will face problems of distortion from partially submerged participants, light artifacts from surface reflection and noise from turbulent water [283, 112]. In this section we link work concerning more research on object tracking and human pose estimation and specifically highlight cases that could inform live shot selection or real-time image correction methods.
@@ 3.3.1 Human Pose Estimation
Human pose estimation (HPE) is a popular paradigm in computer vision research [10, 235, 311]. We see cases of HPE used in cinematic scene analysis [277, 286] and also in live event broadcasting. For example, in sporting events, analyzing a player’s pose may allow us to to forecast a series of entertaining events which could provoke particular cinematographic shots. For sport the major factor that divides research is the environment. We make this delineation prior to reviewing a methods to emphasize that as solutions are heterogeneous [^11] E.g. anatomical poses will vary between sport. [10]. For example, there are special cases where models take-on specific challenges, such as aligning human poses across multiple scenes using 3D pose estimation [277, 249], in a highly dynamic environments [104] or for team sports [28].
Hu [113] looks at real-time football player (single-target) posture estimation for live presentational analysis. To accomplish this, confidence weighted fusion and visual attention is used handle problems with color camouflage and static foreground features, to first identify the target foreground features. Figure13 illustrates pixel-based joint verification for identifying key-point target features, using local binary similarity patterns (lbsp). Then a heat map is generated using a ResNet, a stacked hour glass network and deep-layer aggregation (DLA). The DLA collects features from CNN layers to determine the shared features, i.e. it aggregates features at each layer relating to a verified joint and tries to classify what it is (knee, elbow, etc.).For optimization, the model is updated with adaptive strategies relating to the confidence weight of each pixel and their corresponding weighted fusion sum - based on the joint classification. This methods shows significant improvement in joint and posture detection compared to other state-of-the-art methods [269, 200, 111, 194].
(a) Joint-verification using color features, $v_{t}$, and the local binary similarity pattern, $lbsp_{t}$, of pixel $x$ at frame $t$. $R_{color}$ and $R_{lbsp}$ are predefined distance thresholds and $\{v_{1},v_{2},...,v_{n}\}$ and $\{lbsp_{1},lbsp_{2},...,lbsp_{m}\}$ are a set of known color and binary similarity features. The distance function between the pixel’s color feature, $lbsp$ feature and features in their respective sets, is evaluated w.r.t the thresholds and the binary result is fused using a logical AND operation to determine where a pixel belongs to a joint.
@@
Figure 13: Per-pixel binary joint verification used for identifying pixels for joint classification
In contrast, Bridgeman et al [28] looks at tracking and analysing multiple targets simultaneously. This method builds upon existing research by Cao et al [37]. The goal is to construct a 3D (skeletal) pose model using multiple 2D views, while tackling prior issues of long processing time and reliance on appearance-based queues for initiating feature detection. This is achieved by presenting a greedy approach to multi-view detection. The approach has three steps: (1) 2D Pose Error Correction is done by flipping incorrectly oriented body parts and dividing by body part, (2) Per-frame 2D pose association determines a consistent label for each body part across multiple views, found greedily by selecting the best pose from a weighted rank, and (3) 3D skeleton tracking uses the 2D labelled poses to generate a 3D skeleton for each individual. The novelty in this paper is it’s use of multi-view HPE to verify joint placements in 3D.